# Gravity versus astrophysics in black hole images and photon rings: equatorial emissions and spherically symmetric spacetimes

### v1

@@ C.6 $a_{0}$ and $b_{0}$
Figure 17: Positions of the $n=1$ peaks for all the possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with non-vanishing parameter $b_{1}=0.1$ (blue) and a RZ black hole with non-vanishing parameters $a_{0},\,b_{0}=10^{-4}$ (yellow) or $a_{0},\,b_{0}=0.1$ (red). Critical curves are represented as dashed lines: the green and blue ones are superposed while the yellow one is slightly above them.
If PPN constraints are satisfied, fixing $a_{0}$ and $b_{0}$ to zero, as we have chosen to do in this article, is legit and useful to analyse the impact of more influential parameters. In fact, as can be seen from the orange bands of Figure C.5, comprising the positions of the intensity peaks for all the emission profiles of Table 2, a spacetime defined by the maximum allowed values of $a_{0}$ and $b_{0}$ from stellar system observations, that is, $a_{0}=b_{0}=10^{-4}$ with all other parameters vanishing, produces indiscernible differences from the Schwarzschild green case. If PPN constraints are not satisfied around M87*, which is possible in theories of gravity without a Birkhoff-like uniqueness theorem, the $a_{0}$ parameter is a relevant deviation to consider as it leads to observables comparable to those produced by $a_{1}$, see Figure C.5.
@@ C.7 $b_{1}$
We repeat here the same study on the peaks positions for small deviations in the parameter $b_{1}$. Particularly, we choose $b_{1}=0.1$, similarly to what have been done for $\epsilon$ or $a_{1}$, and we fix all the other metric parameters to their Schwarzschild value. The results of this spacetime configuration are contained in the blue bands of Figure C.5, which does not significantly stand out from the green Schwarzschild band. In other words, small geometrical deviations in $b_{1}$ are degenerate with the astrophysical uncertainty both for the primary image and for the first-lensed photon ring, because $b_{i}$ parameters, with $i\in\mathbb{N}$, do not affect any of the spacetime properties related to the key image features, as explained in Section 6. Hence, this justifies our decision to ignore its effects in the main exploration of this article. We note that even for the extreme value $b_{1}=1$, the degeneracy is never broken.
Figure 18: Widths of the $n=0$ (left panels) and $n=1$ (right panels) images for all the possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=0.1$ (pink) and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). The velocity of the disc is taken to be Keplerian for the upper panels, radial for the central panels and mixed for the lower panels.
Figure 19: Positions of peaks and widths of the $n=0$ (odd panels) and $n=1$ (even panels) profiles for all the possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=1.0$ (pink) and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.5$ (violet). The velocity of the disc is taken to be Keplerian for first row, radial for second row and mixed for the third row. Critical curves are represented as dashed lines.
@@ C.8 $a_{1}$ and $\epsilon$
For small geometric deviations in $\epsilon$ and $a_{1}$, the width of the first photon ring is not an observable permitting to detect deviations from the Schwarzschild geometry, as can be seen in Figure C.6. However, for higher values of these same parameters, the width can become a complementary probe to the peak positions, although the latter remains a more robust probe. These statements can be visualised in Figures C.6 and C.7 where we represent the two extreme positive values allowed by conditions (3.2), namely $a_{1}|_{\,\epsilon=0}=1$ and $\epsilon\,|_{\,a_{1}=0}=0.5$. As already discussed in the main text, larger deviations require less stringent $M/D$ precisions to be observed. We remark also that these extreme deviations lead to more striking differences in the extent, but not the shape, of the bands which is a feature linked to the size of the corresponding lensing bands, which, despite not being observable, must contain the photon rings entirely (Cárdenas-Avendaño & Held 2024). Even for extreme RZ parametric deviations, the astrophysical-driven degeneracy on the image asymmetry and images size ratios introduced in Section 8.2 is never broken. Thus, relative measurements can never be used to distinguish a Schwarzschild black hole from a Rezzolla-Zhidenko spacetime.

### v2

@@ C.6 $a_{0}$ and $b_{0}$
Figure 17: Positions of the $n=1$ peaks for all the possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with non-vanishing parameter $b_{1}=0.1$ (blue), and a RZ black hole with non-vanishing parameters $a_{0},\,b_{0}=10^{-4}$ (yellow) or $a_{0},\,b_{0}=0.1$ (red). Dashed lines represent critical curves: the green and blue ones are superposed while the yellow one is slightly above them.
If PPN constraints are satisfied, fixing $a_{0}$ and $b_{0}$ to zero, as we chose to do in this article, is legit and useful to analyse the impact of more influential parameters. In fact, as can be seen from the orange bands of Figure C.5, comprising the positions of the intensity peaks for all the emission profiles of Table 2, a space-time defined by the maximum allowed values of $a_{0}$ and $b_{0}$ from Solar System observations, that is, $a_{0}=b_{0}=10^{-4}$ with all other parameters vanishing, produces indiscernible differences from the Schwarzschild green case. If PPN constraints are not satisfied around M87*, which is possible in theories of gravity without a Birkhoff-like uniqueness theorem, the $a_{0}$ parameter is a relevant deviation to consider as it leads to observables comparable to those produced by $a_{1}$ (see Figure C.5).
@@ C.7 $b_{1}$
We repeated the same study on the peaks positions for small deviations in the parameter $b_{1}$. Particularly, we chose $b_{1}=0.1$, similarly to what was done for $\epsilon$ or $a_{1}$, and we fixed all the other metric parameters to their Schwarzschild value. The results of this space-time configuration are contained in the blue band of Figure C.5, which does not significantly stand out from the green Schwarzschild band. In other words, small geometrical deviations in $b_{1}$ are degenerate with the astrophysical uncertainty both for the primary image and for the first-lensed photon ring, because $b_{i}$ parameters, with $i\in\mathbb{N}$, do not affect any of the space-time properties related to the key image features, as explained in Section 6. Hence, this justifies our decision to ignore its effects in the main exploration of this article. We note that even for the extreme value $b_{1}=1$, the degeneracy is never broken.
@@ C.8 $a_{1}$ and $\epsilon$
Figure 18: Widths of the $n=0$ (left panels) and $n=1$ (right panels) images for all the possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=0.1$ (pink), and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). The velocity of the disc is taken to be Keplerian for the upper panels, radial for the central panels, and mixed for the lower panels.
Figure 19: Positions of peaks and widths of the $n=0$ (odd panels) and $n=1$ (even panels) profiles for all the possible emission configurations in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=1.0$ (pink), and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.5$ (violet). The velocity of the disc is taken to be Keplerian for first row, radial for second row, and mixed for the third row. Critical curves are represented as dashed lines.
For small geometric deviations in $\epsilon$ and $a_{1}$, the width of the first photon ring is not an observable that permits detection of deviations from the Schwarzschild geometry, as can be seen in Figure C.6. However, for higher values of these same parameters, the width can become a complementary probe to the peak positions, although the latter remain a more robust probe. These statements can be visualised in Figures C.6 and C.7 where we represented the two extreme positive values allowed by conditions (3.2), namely $a_{1}|_{\,\epsilon=0}=1$ and $\epsilon\,|_{\,a_{1}=0}=0.5$. As already discussed in the main text, larger deviations require less stringent $M/D$ precisions to be observed. We remark also that these extreme deviations lead to more striking differences in the extent, but not the shape, of the bands. This is a feature linked to the size of the corresponding lensing bands, which, despite not being observable, must contain the photon rings entirely (Cárdenas-Avendaño & Held 2024). Even for extreme RZ parametric deviations, the astrophysical-driven degeneracy on the image asymmetry and images size ratios introduced in Section 8.2 is never broken. Thus, relative measurements can never be used to distinguish a Schwarzschild black hole from a RZ space-time.